ADNAN MUHAMMED

Grade 12,

a body has mass 40kg and moving with 40m/s velocity what kinetic energy

a body has mass 40kg and moving with 40m/s velocity what kinetic energy

Grade:6

11 Answers

mayank
10 years ago

1/2 * 40 *402

vinit pathak
10 years ago

Use k =1/2mv2

Then u will get ans 32000

sumanth
10 years ago

we know that formula for kinetic enrgy ,K.E=1/2mv2

here mass =40kgs

      velocity=40m/s

     by formula=0.5*40*40*40

                    =32000

naniuppati
10 years ago

given,mass=40kg

velocity=40m/sec

from the formula of kinetic energy    1/2mv^2 =1/2*40*40^2=32000kg.m/s^2

vijay
10 years ago

ans;32000,kinetic energy=1/2mv2=1/2*40*40*40

Amit Sinha
10 years ago
kinetic energy = 1/2*m*sq.v 1/2*40*40=3200joule
sai
10 years ago
K.E=1/2MVsquare. ans:1/2(40)(1600)=20(1600)=3200j
Aditya Shet
10 years ago
K.E=0.5xmxv^2=0.5x40x40^2=3200J
sivakrishna
10 years ago
Given M=40kg V=40m/s According to the equation K=1/2 m v^2 =1/2(40)(40)^2 =64000/2 =32000
BIRADHER SHREYA
10 years ago
32000m/s
THEJA
9 years ago
kinetic enrgy ,K.E=1/2mv2 here mass =40kgs velocity=40m/s by formula=0.5*40*40*40 =32000